OpenStudy (anonymous):

If f(4) = 8, g(4) = 2, f(-4) = 1, g(-4) = 3, !'(4) = 10, g'(4) = 12, f'(-4) = 6, and g'( -4) = -2, find H'(4) where H(x) = g( f(x)- 3x).

OpenStudy (anonymous):

H'(x)= g'(f(x)-3x)(f'(x)-3) H'(4)= g'(f(4)-12)(f'(4)-3) = g'(8-12)(10-3) =g'(-4)(7) =-2*7 =-14.